OpenStudy (anonymous):

multiply and simplify 4x1/2

OpenStudy (anonymous):

is the answer 2/2?

OpenStudy (chosenmatt):

no it's 2

OpenStudy (anonymous):

The answer is two because you multiply 4/1 by 1/2 and you get 4/2 or 2

OpenStudy (anonymous):

\[4 * 1 \div 2 ?\]

OpenStudy (anonymous):

the 1/2 is in fraction

OpenStudy (anonymous):

\[4 * \frac{ 1 }{ 2 } \]

OpenStudy (chosenmatt):

click the link

OpenStudy (anonymous):

i got it wrong the answer is 1 :(

OpenStudy (chosenmatt):

ouch

OpenStudy (chosenmatt):

srry

OpenStudy (anonymous):

what about 45x1/9

OpenStudy (chosenmatt):

answer is 5

OpenStudy (anonymous):

can i put 5/1?

OpenStudy (chosenmatt):

well here's the work (45×1)÷9=5

OpenStudy (anonymous):

what about 4/3 x 1/4

OpenStudy (anonymous):

has to be simplify and multiply

OpenStudy (chosenmatt):

((1÷4)×4)÷3= 0.333333333

OpenStudy (chosenmatt):

s and m

OpenStudy (chosenmatt):

ok bye